# The Impact of Prompt Programming on Function-Level Code Generation

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly used by software engineers for code generation. However, limitations of LLMs such as irrelevant or incorrect code have highlighted the need for prompt programming (or prompt engineering) where engineers apply specific prompt techniques (e.g., chain-of-thought or input-output examples) to improve the generated code. Despite this, the impact of different prompt techniques — and their combinations — on code generation remains underexplored. In this study, we introduce CodePromptEval, a dataset of 7072 prompts designed to evaluate five prompt techniques (few-shot, persona, chain-of-thought, function signature, list of packages) and their effect on the correctness, similarity, and quality of complete functions generated by three LLMs (GPT-4o, Llama3, and Mistral). Our findings show that while certain prompt techniques significantly influence the generated code, combining multiple techniques does not necessarily improve the outcome. Additionally, we observed a trade-off between correctness and quality when using prompt techniques. Our dataset and replication package enable future research on improving LLM-generated code and evaluating new prompt techniques.
@@ I Introduction
These limitations are inherent to the design of LLMs, and are unlikely to ”resolve themselves” entirely with future model generations. Therefore, researchers started proposing ways to mitigate these limitations by adapting how users interact with the LLMs. The interactions typically start with a natural language prompt that specifies what the LLM is expected to output. To ensure that the LLM generates accurate, relevant, and high-quality outputs, users employ a structured approach to construct prompts, which is often referred to as prompt programming.
@@
However, such prompt techniques were evaluated based on the output accuracy for natural language generation tasks [9, 8] and are not well-studied for code generation, more specifically, function synthesis (generating function-level code), which is one of the most common use cases among software engineers [3]. Furthermore, evaluating the accuracy of code generation is not sufficient, since other aspects of the code are important for software engineers, such as maintainability and adherence to best practices. Prompt techniques can also be combined [6], but to the best of our knowledge, no work evaluates the impact of multiple combinations of prompt techniques in one prompt. For instance, whether applying a certain prompt technique can cancel out, hinder, or even enhance the impact of an existing prompt technique in the prompt.
@@ II Related work
For code-related tasks, prompt techniques were shown to have a positive effect on code generation in the domain of education [22]. Furthermore, researchers proposed ways and contextual information as prompt techniques to apply to the prompt and enhance code-related tasks [23, 8], e.g., incorporating dataflow information to improve code summarization [8]. Other prompt techniques used by Dong et al [24] included self-collaboration, where the LLM is prompted several times to take different personas e.g., first as a requirements engineer, then a software developer, then a tester, and only then return a code that resulted from the “collaboration” among the three personas. In our study, we focus on three common prompt techniques, namely, few-shot learning, automatic chain-of-thought [25], and persona [26], as well as propose two pieces of contextual information as additional prompt techniques that are easily accessed by the developers, i.e., the imported packages and the signature of the function.
Figure 1: The process we follow to evaluate the code generated using different prompts by different LLMs.
@@ III-A Prompt technique combinations
  • Few-shot learning can be achieved by providing shots (or examples) to an LLM in order to enable learning new examples without the need to fine-tune the LLM [19]. We use two input-output examples explained in natural language. We do not consider a varying number of shots.
  • Automatic Chain-of-Thought (CoT) allows the LLM to break down the prompt by asking it to ”think” step by step and following the steps when solving the problem [25].
@@
  • Packages is a list of libraries and files that exist in the environment in which the code runs. This includes local packages and external libraries.
@@ III-B CodePromptEval
The next step was to prepare prompt templates by defining how each prompt technique will be implemented and mapping relevant information to prompt techniques. In particular, for each datapoint, we extract the signature of the function and the list of used packages (represented as imports at the beginning of the class). For chain-of-thought, we adapted the template recommended by Zhuosheng et al. [25]. To construct the persona, we defined a persona description of a software developer who follows best coding practices for maintainability. To implement the few-shot prompt technique, the first three authors of this paper manually constructed two input-output examples for each prompt following the template “If the input is X, then the output is Y”. We also create corresponding tests to ensure that the input and output are correct.
@@
We then map each variation from Table I to the relevant information and templates for prompt techniques (e.g., imported libraries for packages), then we combine them with the 221 prompts from CoderEval, leading to 7072 concrete prompts (221 prompts times 32 variations). Our dataset, the virtual environments, and the few-shot examples and tests are provided in our replication package [35].
@@ III-C Code Generation
We focus on LLMs with a decoder-only transformer architecture, which is at the time of writing the preferred architecture to use in code generation tasks [36]. Therefore, we select the following LLMs for our study: GPT-4o, Llama3-70B-Instruct, and Mistral-Small-Instruct-2409 (22B). We also collect data for two previous-generation LLMs (GPT-3.5-turbo and Llama2-7B-Instruct), but omit discussing the results for these older models for reasons of brevity in this paper. The collected data for these models is still available in our replication package [35].
We run all 7072 prompts on the selected LLMs with 0.2 temperature, which has been commonly used for code generation tasks [27, 37]. For the API-based GPT models, we send requests to the external API and store the responses. We host the remaining models on the Alvis cluster, a NAISS resource (National Academic Infrastructure for Supercomputing in Sweden) dedicated to Artificial Intelligence and Machine Learning research [^4] https://www.c3se.chalmers.se/about/Alvis/ using models downloaded from Huggingface [^5] https://huggingface.co. Running the self-hosted LLMs on Alvis required around 1600 GPU hours using Nvidia A100 GPUs. For the GPT models, we use the OpenAI API, which is billed based on the tokens that are processed. To run GPT-4o and GPT-3-turbo on all the prompts in CodePromptEval, we provide around 1.1 million input tokens and generate approximately 3.7 million output tokens.
@@ III-D Evaluating the LLM-generated functions
After generating 7072 code solutions, we evaluate them based on three main aspects following our research questions (correctness, similarity, and quality). We use different tests to measure statistically significant differences for the measures below, hence we detail the choice of statistical methods in their corresponding results sections.
@@
To evaluate their correctness, we run the generated functions against their corresponding tests in CoderEval. There are two types of tests in CoderEval: Python unit tests, and a main function with different statements and conditions that set a boolean variable isT (is True) to False when at least one of the conditions does not hold. To ensure consistency and instrumentation of our experiment, ensure that an AssertionError is thrown when needed, by adding an assert statement at the end of tests in the form of a main function assert isT. Furthermore, as some of the LLM-generated functions can be erroneous and get stuck in an infinite loop, we wrap the tests with error-handling constructs (a try/except) and set a timeout of 60 seconds per function. Then we collect the test results and the error messages when applicable.
@@
To assess the LLM-generated code’s similarity to the ground truth obtained from CoderEval (human-written functions), we measure the CodeBLEU score [29]. CodeBLEU combines four dimensions of similarity to capture alignment in the syntax, semantics, and data flow between the generated functions and the ground truth. The different dimensions are N-gram match (BLEU), weighted N-gram matches (weighted-BLEU), Abstract Syntax Tree (AST) match, and data flow match. For more specific insights, we also measure the syntactic similarity (coding style and variable naming) using BLEU and weighted-BLEU, and semantic similarity (i.e., structural and algorithmic design) using AST match.
@@ IV CodePromptEval Overview
In this section, we provide an overview of the aggregate results from running three LLMs (GPT-4o, Llama-3, and Mistral) on the CodePromptEval dataset. This section answers RQ1 in our study.
@@
TABLE II: Overview of passed and failed functions per LLM. The total number of functions per LLM is 7072.
LLM # Passed Functions # Failed Functions
GPT-4o 3707 (52.42%) 3365 (47.58%)
Llama3-70B-Instruct 3564 (50.40%) 3508 (49.60%)
Mistral-22B-Instruct 3335 (47.16%) 3737 (52.84%)
@@
A high-level results summary is shown in Table II. There are a total of 7072 generation tasks in the dataset. All three models are able to solve (generate functions that pass all tests) approximately half of the tasks. Mistral performs worst in our study, solving 3335 (47.16%) of tasks, and GPT-4o does best solving 3707 (52.42%), outperforming the worst model by approximately 5 percentage points.
@@
To get a better idea of whether these results are impacted by the code level of the function. We use the code levels defined by Yu et al. [10] that are based on the nature of dependencies of the function. The code levels are: self-contained (no need to import), standard library runnable (no need to install), public library runnable (uses libraries available on PyPI), class runnable (uses code outside the function, but within the class), file runnable (uses code outside the class, but within the file), and project runnable (uses code in other files). Code levels provide a rough indication of the “difficulty” of a generation task, based on what kind of dependencies the LLM needs to correctly incorporate.
@@
Figure 3: Passed and failed functions per LLM for each code level. The total number of functions per LLM is 7072.
@@
We looked into the code levels that passing and failing functions belong to (see Figure 3). Unsurprisingly, the fail rate for all models increases as tasks get more difficult (i.e., by construction, class runnable tasks tend to be substantially more challenging than self-contained ones, and all models struggle much more with solving them correctly). Pass rates for the easiest type of task (standard library runnable) are close to 90% for all models, going down to as low as 31% to 41% for the most challenging tasks (class runnable). We observe that, overall, all three models perform comparably on most code levels, with the exception of self-contained tests (where GPT-4o outperforms the other models by a larger margin of 10 to 12 percentage points). This difference explains most of the slightly higher overall performance of GPT-4o. We have also confirmed these differences using Chi-square statistical test and Cohen’s $\omega$ as effect size ($\omega=0.34$ :- Medium effect) [40].
@@
Figure 4: Frequency of error types among failing tests for functions generated by GPT-4o, Llama3-70B, and Mistral-22B.
@@
Finally, we report what error types led to the failing tests shown in Table II and Figure  3. We report the error types based on the Python exception that is first thrown when running the tests. The results of the error types are visualized in Figure 4. The most common error type for failed tests across the LLMs is AssertionError, indicating that the LLM generated a Python function that did not exhibit precisely the expected functionality (as defined through unseen tests). However, are also frequently encountered such as TypeErrors (operation is performed on a value of an inappropriate type, indicating that the LLM misjudged the runtime type of a Python object), AttributeErrors (invalid attribute reference is made), and ImportErrors (a faulty import of a module or object). Other errors, such as NameErrors or IndexErrors, exist but are rare. While there are differences between the LLMs, they are relatively minor and not systematic. The most notable difference is that Mistral tends to generate functions leading to an ImportError or NameError more frequently than the other LLMs, whereas AttributeErrors are less frequent in Mistral-generated code.
@@
Key Findings (RQ1): Overall, we observed that GPT-4o minorly outperforms the other LLMs in the study. However, in general, results are consistent between current-generation LLMs. Depending on task difficulty, all LLMs can solve between 31% and 89% of tasks. Assertion and TypeErrors are the most common cause of failed tests.

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly used by software engineers for code generation. However, limitations of LLMs such as irrelevant or incorrect code have highlighted the need for prompt programming (or prompt engineering) where engineers apply specific prompt techniques (e.g., chain-of-thought or input-output examples) to improve the generated code. While some prompt techniques have been studied, the impact of different techniques — and their interactions — on code generation is still not fully understood. In this study, we introduce CodePromptEval, a dataset of 7072 prompts designed to evaluate five prompt techniques (few-shot, persona, chain-of-thought, function signature, list of packages) and their effect on the correctness, similarity, and quality of complete functions generated by three LLMs (GPT-4o, Llama3, and Mistral). Our findings show that while certain prompt techniques significantly influence the generated code, combining multiple techniques does not necessarily improve the outcome. Additionally, we observed a trade-off between correctness and quality when using prompt techniques. Our dataset and replication package enable future research on improving LLM-generated code and evaluating new prompt techniques.
Index Terms:
Large Language Models, Prompt Programming, Code Generation.
@@ I Introduction
These limitations are inherent to the design of LLMs, and are unlikely to ”resolve themselves” entirely with future model generations. Therefore, researchers started proposing ways to mitigate these limitations by adapting how users interact with the LLMs. The interactions typically start with a natural language prompt that specifies what the LLM is expected to output. To ensure that LLM generates accurate, relevant, and high-quality outputs, users employ a structured approach to construct prompts, which is known as prompt programming.
@@
However, such prompt techniques were evaluated based on the output accuracy for natural language generation tasks [9, 8] and are not well-studied for code generation, more specifically, function synthesis (generating function-level code), which is one of the most common use cases among software engineers [3]. Furthermore, evaluating the accuracy of code generation is not sufficient, since other aspects of the code are important for software engineers, such as maintainability and adherence to best practices. Prompt techniques can also be combined [6], but to the best of our knowledge, no work evaluates the impact of multiple interacting prompt techniques in one prompt. For instance, whether applying a certain prompt technique can cancel out, hinder, or even enhance the impact of an existing prompt technique in the prompt.
@@ II Related work
For code-related tasks, prompt techniques were shown to have a positive effect on code generation in the domain of education [22]. Furthermore, researchers proposed ways and contextual information as prompt techniques to apply to the prompt and enhance code-related tasks [23, 8], e.g., incorporating dataflow information to improve code summarization [8]. Other prompt techniques used by Dong et al [24] included self-collaboration, where the LLM is prompted several times to take different personas e.g., first as a requirements engineer, then a software developer, then a tester, and only then return a code that resulted from the “collaboration” among the three personas. Fagadau et al. [25] examined how individual prompt features (active voice and edge cases) affect code generation. Their experiments, which combined different prompt features, showed that most had little impact on the resulting code. In our study, we focus on three common prompt techniques, namely, few-shot learning, chain-of-thought [26], and persona [27], as well as propose two pieces of contextual information as additional prompt techniques that are easily accessed by the developers, i.e., the imported packages and the signature of the function. In addition, we investigate not only the impact of individual prompt techniques but also their interaction effects on code generation, for example, whether the effect of combining few-shot and persona arises from their interaction or from one technique alone.
Figure 1: The process we follow to evaluate the code generated using different prompts by different LLMs (per run).
@@ III-A Prompt technique combinations
  • Few-shot learning can be achieved by providing shots (or examples) to an LLM in order to enable learning new examples without the need to fine-tune the LLM [19]. Typically, the examples describe the structure of the input and output. In code generation, such pattern will result in each example being composed of a generation task in natural language as an input, and a complete function as an output. We found this to be impractical from a user perspective, and poses a challenge of what generation tasks to choose in terms of the prompt design. Therefore, we follow an adapted pattern of few-shot prompting for code generation tasks also used in previous work [28, 34], where each example consists of a possible input of the function and its corresponding output. We use two input-output examples explained in natural language. We do not consider a varying number of shots.
  • Chain-of-Thought (CoT) allows the LLM to break down the prompt by asking it to “think” step by step before solving the problem. This technique is used to prompt the LLM to perform explicit reasoning [35]. We apply Zero-shot-CoT [26] to isolate the impact of CoT from the few-shot prompt technique. In Zero-shot-CoT the steps that the LLM can follow are not explicitly mentioned in the prompt; rather, the model is expected to generate and follow its own reasoning process autonomously.
@@
  • Packages is a list of libraries and files that exist in the environment in which the code runs. This includes local packages and external libraries. The packages used by a function are extracted by parsing all import statements in the Python file to collect the names of the modules and external libraries on which the function depends.
Previous work indicates that Signature and Packages are not necessarily used when prompting LLMs for code generation by developers [3]. While Signature is often a part of the prompt in code completion benchmarks [38], most code generation benchmarks construct prompts based on documentation (e.g., docstrings) without the signature  [10]. Therefore, in this study, we present them as prompt techniques for code generation tasks as they might provide additional context that can guide the model toward more relevant code generation (e.g., types inferred by the parameter names).
@@ III-B CodePromptEval
The next step was to prepare prompt templates by defining how each prompt technique will be implemented and mapping relevant information to prompt techniques. In particular, for each datapoint, we extract the signature of the function and the list of used packages (represented as imports at the beginning of the class). For chain-of-thought, we adapted the template recommended by Zhuosheng et al. [26]. To construct the persona, we defined a persona description of a software developer who follows best coding practices for maintainability. To implement the few-shot prompt technique, the first three authors of this paper manually constructed two input-output examples for each prompt following the template “If the input is X, then the output is Y”. We also create corresponding tests to ensure that the input and output are correct. The examples were created based on the goal of covering both a typical (mainline) case and an edge case to ensure that the examples capture a range of expected behavior.
@@
We then map each variation from Table I to the relevant information and templates for prompt techniques (e.g., imported libraries for packages), then we combine them with the 221 prompts from CoderEval, leading to CodePromptEval with 7072 concrete prompts (221 prompts times 32 variations) and their corresponding Python functions as ground truth.
The functions include both domain-specific implementations and commonly used utility logic extracted from GitHub repositories, rather than standard textbook algorithms such as sorting or searching. Table II describes the functions’ length and cyclomatic complexity. Moreover, the 221 functions fall into six code dependency levels: 33 self-contained (does not use packages outside the function scope), 25 standard library runnable (uses libraries available as part of Python standard library), 19 public library runnable (uses libraries available on PyPI), 54 class runnable (uses code outside the function, but within the class), 67 file runnable (uses code outside the class, but within the file), and 23 project runnable (uses code in other files). Our dataset, the virtual environments, and the few-shot examples and tests are provided in our replication package [39].
TABLE II: Function Statistics in CodePromptEval
Metric Min Max Mean Std Dev
Number of Variables 0 32 2.52 4.35
Number of Parameters 0 7 1.70 1.30
Lines of Code 3 564 32.14 55.82
Block Depth 1 9 2.58 1.63
Cyclomatic Complexity 1 29 4.57 4.66
@@ III-C Code Generation
We focus on LLMs with a decoder-only transformer architecture, which is at the time of writing the preferred architecture to use in code generation tasks [40]. Therefore, we select the following LLMs for our study: GPT-4o, Llama3-70B-Instruct, and Mistral-Small-Instruct-2409 (22B). We also collect data for two previous-generation LLMs (GPT-3.5-turbo and Llama2-7B-Instruct), but omit discussing the results for these older models for reasons of brevity in this paper. In general, the results for these older models showed lower passing rates and higher complexities. However, the overall impact of the techniques remained consistent with the findings we report for the studied LLMs, albeit with varying significance levels. The collected data for these models is still available in our replication package [39].
We run all 7072 prompts on the selected LLMs three times to account for the non-deterministic nature of LLMs. For all LLMs, we set the temperature to 0.2, which has been commonly used for code generation tasks [28, 41], and complies with the recommendations for our correctness measure [28]. For the API-based GPT models, we send requests to the external API and store the responses. We host the remaining models on the Alvis cluster, a NAISS resource (National Academic Infrastructure for Supercomputing in Sweden) dedicated to Artificial Intelligence and Machine Learning research [^4] https://www.c3se.chalmers.se/about/Alvis/ using models downloaded from Huggingface [^5] https://huggingface.co. Running the self-hosted LLMs on Alvis required around 2800 GPU hours using Nvidia A100 GPUs. For the GPT models, we use the OpenAI API, which is billed based on the tokens that are processed. To run GPT-4o and GPT-3-turbo on all the prompts in CodePromptEval, we provide around 2.45 million input tokens and generate approximately 7.66 million output tokens.
@@ III-D Evaluating the LLM-generated functions
After generating 7072 code solutions three times, we evaluate them based on three main aspects following our research questions (correctness, similarity, and quality). We use different tests to measure statistically significant differences for the measures below, hence we detail the choice of statistical methods in their corresponding results sections.
@@
To evaluate their correctness, we run the generated functions against their corresponding tests in CoderEval. When running the functions, we replace the generated function name with the originally expected one to ensure compliance with our test cases.
There are two types of tests in CoderEval: Python unit tests, and a main function with different statements and conditions that set a boolean variable isT (is True) to False when at least one of the conditions does not hold. To ensure consistency and instrumentation of our experiment, ensure that an AssertionError is thrown when needed, by adding an assert statement at the end of tests in the form of a main function assert isT. Furthermore, as some of the LLM-generated functions can be erroneous and get stuck in an infinite loop, we wrap the tests with error-handling constructs (a try/except) and set a timeout of 60 seconds per function. Then we collect the test results and the error messages when applicable.
We distinguish syntactic correctness and semantic correctness of the function. The Python function is syntactically correct if its syntax is valid and the function is runnable. Semantically correct functions are functions that pass their corresponding tests. When a function is both syntactically and semantically correct, then it is labeled as plausibly correct [42]. In the remainder of the paper, we use correctness as a short-hand for plausibly correct.
@@
To assess the LLM-generated code’s similarity to the ground truth obtained from CoderEval (human-written functions), we measure the CrystalBLEU score [43]. CrystalBLEU combines four n-gram measures where $n=1,2,3,4$ while accounting for “trivial grams” that are shared across all functions, such as Python keywords. The combined n-grams that are used as a metric for syntactic similarity are then used as a proxy to estimate the semantic similarity.
@@ IV CodePromptEval Overview
In this section, we provide an overview of the aggregate results from running three LLMs (GPT-4o, Llama-3, and Mistral) three times on the CodePromptEval dataset. This section answers RQ1 in our study.
@@
TABLE III: Overview of passed and failed functions per LLM. We also show the breakdown of failures types (total = 7072 functions, averaged over 3 runs)
Results GPT-4o Llama3-70B Mistral-22B
Passed 3691 $\pm$ 12 (52.2%) 3575 $\pm$ 11 (50.5%) 3318 $\pm$ 13 (46.9%)
Failed 3381 $\pm$ 12 (47.8%) 3497 $\pm$ 11 (49.5%) 3754 $\pm$ 13 (53.1%)
- Syntactic 27 $\pm$ 1 (0.4%) 59 $\pm$ 4 (0.9%) 103 $\pm$ 3 (1.5%)
- Semantic 1303 $\pm$ 23 (18.4%) 1415 $\pm$ 15 (20.0%) 1182 $\pm$ 18 (16.7%)
- Operational 2051 $\pm$ 21 (29.0%) 2023 $\pm$ 6 (28.6%) 2468 $\pm$ 16 (34.9%)
@@
A high-level results summary is shown in Table III. There are a total of 7072 generation tasks in the dataset. All three models are able to solve (generate functions that pass all tests) approximately half of the tasks. Mistral performs worst in our study, solving on average 3318 (46.9%) of tasks, and GPT-4o does best solving 3691 (52.2%), outperforming the worst model by approximately 5 percentage points.
@@
To get a better idea of whether these results are impacted by the code level of the function, we use the code levels defined by Yu et al. [10] that are based on the nature of dependencies of the function. The code levels are: self-contained, standard library runnable, public library runnable, class runnable, file runnable, and project runnable. Code levels provide a rough indication of the “difficulty” of a generation task, based on what kind of dependencies the LLM needs to correctly incorporate.
@@
Figure 3: Passed and failed functions per LLM for each code level across three runs. The total number of functions per LLM in a single run is 7072.
@@
We looked into the code levels that passing and failing functions belong to (see Figure 3). Unsurprisingly, the fail rate for all models increases as tasks get more difficult (i.e., by construction, class runnable tasks tend to be substantially more challenging than self-contained ones, and all models struggle much more with solving them correctly). Pass rates for the easiest type of task (standard library runnable) are close to 90% for all models, going down to as low as 31% to 41% for the most challenging tasks (class runnable). We observe that, overall, all three models perform comparably on most code levels, with the exception of self-contained tests (where GPT-4o outperforms the other models by a larger margin of 10 to 13 percentage points). This difference explains most of the slightly higher overall performance of GPT-4o. We also confirmed these differences using Chi-square test, which assesses the association between categorical variables (code level and pass/fail outcome) resulting in p-value $<$ 0.0001, and Cramér’s $\phi$ as an effect size for the relationship between the two nominal variables ($\phi=0.34$ :- medium effect).
@@
Figure 4: Percentages of error types occurring among failing tests for functions generated by GPT-4o, Llama3-70B, and Mistral-22B across three runs.
@@
Finally, we report what errors led to the failing tests shown in Table III and Figure  3. We report the error types based on the Python exception that is first thrown when running the tests. The results of the error types are visualized in Figure 4. The most common error type for failed tests across the LLMs is AssertionError, indicating that the LLM generated a Python function that did not exhibit precisely the expected functionality (as defined through unseen tests). However, are also frequently encountered such as TypeErrors (operation is performed on a value of an inappropriate type, indicating that the LLM misjudged the runtime type of a Python object), AttributeErrors (invalid attribute reference is made), and ImportErrors (a faulty import of a module or object). Other errors, such as NameErrors or IndexErrors, exist but are rare. While there are differences between the LLMs, they are relatively minor and not systematic. The most notable difference is that Mistral tends to generate functions leading to an ImportError or NameError more frequently than the other LLMs, whereas AttributeErrors are less frequent in Mistral-generated code.
[…]
@@
skipabove=5pt, linecolor=blue, roundcorner=2pt, innerleftmargin=0.3cm,innerrightmargin=0.4cm, linewidth=0.5pt, footnoteinside=false,backgroundcolor=blue!5 {mdframed}Key Findings (RQ1): Overall, we observed that GPT-4o minorly outperforms the other LLMs in the study. However, in general, results are consistent between current-generation LLMs. Depending on task difficulty, all LLMs can solve between 31% and 90% of tasks. Assertion and TypeErrors are the most common cause of failed tests.